INAF queue observing, 2015 Nov 03-Nov 04

Observer: S. Bisogni, R. Carini
Telescope Operator: Steve Allanson
Support Astronomer: Barry Rothberg
Instrument: LUCI, LBC


* 03:10 UT the humidity is 100% and the temperature is at the dew point.

Night duration (12\xB0 el twilight) 19:00-5:10 ( 02:00-12:10 UT) = 10 h 10m

Observing time: 0 hrs 0 m

Weather time loss: 10 h 10 m (100%)

technical time loss 0 m

Data Summary

At the sunset the sky is covered, humidity 86%, very strong wind.
the dome is closed
* 03:10 UT the humidity is 100% and the temperature is at the dew point.

Project_RUN Objects Filter ExpTime Notes
Calibration Saracco        


09:50 begin MODS2 calibration for Saracco program
  • we modify the calibration scripts of the PI for mods2, the example file is grmodscals_r_mods2.cal in the directory /home/modseng/modsScripts/mods2Calib
    the new files are mods2.3.XLSms1.cal and mods2.4.XLSms2.cal
  • 09:55 execMODS2 mods2.3.XLSms1.cal mods2r.20151104.0069-0073 flat+ arc
  • 10:05 execMODS2 mods2.4.XLSms2.cal mods2r.0074-0078 flat+arc
  • 10:20 execMODS2 grpixflats_r_mods2.cal mods2r.0079-0083 flat slitless in red grating
  • 10:30 execMODS2 slitflat_5as_r_mods2.cal mods2r.0084-0086 5′′ Slit flat in Red Grating mode for the standard star
  • 10:50 execMODS2 grlamps_r_mods2.cal mods2r.0087-0089 5'' slit arc in red grating for standard star, mods2r.0087 (ne lamp) is not good
  • 11:00 execMODS2 bias8K.cal mods2r.0090-0094
  • 11:15 execMODS2 bias3K.cal mods2r.0095-0099
  • 11:30 execMODS2 bias1K.cal mods2r.0100-0104
  • 11:40 we rice the time exposure (1 sec) for the Ne lamp mods2r.0105 couple of lines of the spectra are saturated
  • try again with 0.5 sec mods2r.0106 now is fine

Article text.

-- %USERSIG{RobertaCarini - 2015-11-04}%


Topic revision: r9 - 09 Nov 2015, RobertaCarini
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding Foswiki? Send feedback